OpenStudy (marsmcnugget):

Which of the ordered pairs in the form (x, y) is a solution of this equation? Three x minus begin fraction y over four end fraction equals eleven 3x-y/4=11 (3, –8), (4, 4) A. The first is not a solution, but the second is. B. Both are solutions. C. The first is a solution, but the second is not. D. Neither is a solution.

OpenStudy (marsmcnugget):

@raiders88

OpenStudy (marsmcnugget):

@love_to_love_you @luckycoins888 @Adi3 @Rushwr @Elsa213 @toxicsugar22 @ganeshie8 @Jacob902 @VelvetRoux @sweetburger

OpenStudy (anonymous):

@ytrewqmiswi

OpenStudy (anonymous):

@Shadowknight6263

OpenStudy (ytrewqmiswi):

if any ordered pair like (x1,y1) is a solution to that equation then the equation must satisfy when u put the vaue of x as x1 and the value of y as y1.

OpenStudy (anonymous):

Actually Me too want the answer with details @ytrewqmiswi

OpenStudy (ytrewqmiswi):

3x-y/4=11 ok 1st lets check (3,-8) putting x=3 ans y=-8 in the equation we get-> 3(3)-(-8)/4 =11 9+2=11 11=11 since left hand side is coming equal to right hand side (3,-8) is a solution can u check (4,4) urself :)

OpenStudy (marsmcnugget):

is it a b c or d

OpenStudy (ytrewqmiswi):

B

OpenStudy (marsmcnugget):

thanks

OpenStudy (ytrewqmiswi):

np :) if u want jst answer or explanation then jst tell me :)